# A Comprehensive Survey on Uncertainty Quantification for Deep Learning

### v4

@@ 4. Application domains
Biochemistry engineering: Traditional biochemistry discoveries are primarily based on experiments, which are expensive and time-consuming (Bower and Bolouri, 2001). With the rapid advancement of computational science, researchers have turned to simulation and data-driven approaches to aid scientific discovery (Brunton and Kutz, 2022). Simulation methods create mathematical models based on fundamental principles, such as kinetics of microbial growth and product synthesis, fluid dynamics in bioreactors, and design of new proteins and enzymes (Mowbray et al., 2021). However, these models can be computationally intensive and require an in-depth understanding of the biochemical system. In contrast, deep learning-based models use neural networks and a vast amount of biochemistry data to directly extract knowledge and learn complex relationships. However, the common data-driven DNN models are unaware of domain knowledge and face challenges in representing the complexities of biochemical systems, such as the heterogeneity of microbial populations and genetic-environmental variations. This often results in unreliable outputs that require human experts to verify, adding additional labor to the process (Hie et al., 2020). Several factors contribute to the uncertainty of data-driven biochemistry engineering. For example, researchers often have only a partial understanding of the complex biological mechanism and the macromolecules released from the cells and it is difficult to quantify the heterogeneity in the microbial population, due to genetic and environmental variations. For different tasks, these factors may play a different role in the UQ for neural network model. In protein engineering, the protein function prediction (based on protein sequences, protein structures, and protein-protein interactions) is challenging because of the large amount of diversity of protein folds and the lack of a complete understanding of the protein structure (Bradford et al., 2018). For bioreactor engineering, the prediction of bioreactor system performance is affected by the intrinsic interactions between bioreactor operating factors (e.g., pH, temperature, and substrate concentration). An iterative method has been used to enable the propagation of model uncertainty for multi-step ahead predictions (Mowbray et al., 2021).
@@ 5. A Taxonomy of DNN UQ Methodology
In this section, we provide a new taxonomy of existing DNN UQ methods based on the type of uncertainty sources each method can capture. As mentioned before, two main sources of uncertainty exist in the DNN predictions and for different domain applications, different sources may play a major role in affecting the model prediction. Thus it is vital to understand which type of uncertainty source each method can address. In general, we categorize existing works into three branches based on the uncertainty sources as Fig. 4 shows: model uncertainty, data uncertainty, and the combination of the two. We briefly introduce each category as follows:
  • Model uncertainty: In this category, the approaches consider the model uncertainty resulting from the parameters (BNN), architectures (ensemble methods), or sample density (Deep Gaussian process).
  • Data uncertainty: The approaches in this category aim to account for the uncertainty from the inherent randomness or noise in the data. The general idea is to construct a distribution over the prediction. Those approaches are further split into deep discriminative models and deep generative models.
  • The combination of model and data uncertainty: Many approaches aim to capture the total uncertainty in one framework, which is the combination of model and data uncertainty. The straightforward approach is to combine the approaches in data and model uncertainty and form a coherent framework, but introduces more computation and storage demand. Another framework called evidential deep learning overcomes the computational challenge with a single neural network model.
@@ 5.1. Model Uncertainty
This subsection reviews the existing framework for DNN model uncertainty. We categorize the approaches into three subcategories: Bayesian neural network, ensemble models, and sample density aware models. We will introduce each subcategory in detail.
@@ 5.1.1. Bayesian Neural Networks
From a frequentist point of view, there exists a single set of parameters $*\boldsymbol{\boldsymbol{\theta}}*$ that fit the DNN model best, where $\boldsymbol{\theta}*=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathcal{L}(\mathbf{Y},f(\mathbf{X},\boldsymbol{\theta}))$ and $\mathcal{L}$ is the loss function. However, the point estimation of DNN parameters tends to be overfitting and overconfident (Thulasidasan et al., 2019). In contrast, the Bayesian neural network (BNN) imposes a prior on the neural network parameters $p(\boldsymbol{\theta})$ and learns the posterior distribution of these parameters $p(\boldsymbol{\theta}|\mathbf{X},\mathbf{Y})$.
@@
For inference on the new samples $*\boldsymbol{x}*$, we can marginalize out the model parameters by:
@@
However, the approach is analytically intractable and does not have a closed-form solution and an approximation must be made for prediction in BNN. To estimate the posterior of the neural network parameters, various approaches have been proposed to approximate the parameter posterior in a simpler form to solve it in a tractable way. Some approximation methods define a parameterized class of distributions, $\mathcal{Q}$, from which they select an approximation $q_{\phi}(\boldsymbol{\theta})$for the posterior. For example, $\mathcal{Q}$ can be the set of all factorized Gaussian distributions, and $\phi$ is the parameters of mean and diagonal variance. The distribution $q_{\phi}(\boldsymbol{\theta})\in\mathcal{Q}$ is selected according to some optimization criteria to approximate the posterior. Two popular approaches for optimization are variational inference (Blei et al., 2017) and Laplace approximation (Friston et al., 2007). Instead of approximating the posterior in an analytical way, another approach aims to solve this problem by Monto Carlo sampling called Markov Chain Monto Carlo sampling. In the following, we will review existing methodologies in the three subcategories in detail.
@@
One challenge of variational inference lies in the choice of the parameterized class of distribution $q_{\boldsymbol{\phi}}(\boldsymbol{\theta})$. The original method relies on the use of Gaussian approximating distribution with a diagonal covariance matrix (mean-field variational inference) (Hinton and Van Camp, 1993; Posch et al., 2019). This method leads to a straightforward lower bound for optimization, but the approximation capability is limiting. To capture the posterior correlations between parameters, the second category of approaches extends the diagonal covariance to general covariance matrix while still leading to a tractable algorithm by maximizing the above ELBO (Posch and Pilz, 2020). However, the full covariance matrix not only increases the number of trainable parameters but also introduce substantial memory and computational cost for DNN. To reduce the computation, the third category of approaches aim to simplify the covariance matrix structure with certain assumption. Some approaches assume independence among layers, resulting in block-diagonal structure covariance matrix (Sun et al., 2017; Zhang et al., 2018). Others find that for a variety of deep BNN trained using Gaussian variational inference, the posterior consistently exhibits strong low-rank structure after convergence (Swiatkowski et al., 2020), they propose to decompose the dense covariance matrix into a low-rank factorization to simplify the computation. Additionally, (Mishkin et al., 2018) assumes the covariance matrix taking the form of “diagonal plus low-rank” structure for a more flexible and faster approximation. Another line of work introduces sparse uncertainty structure via a hierarchical posterior or employing normalizing flow with low-dimensional auxiliary variables (Ritter et al., 2021; Louizos and Welling, 2017) to reduce the computation.
The disadvantage of variational Gaussian approximation for DNN parameters is that to capture the full correlations among model latent weights, it requires large amount of variational parameters to be optimized, which scales quadratically with the number of latent weights in the model. Existing works aim to simplify the covariance structure with certain assumption while still capturing the correlation between neural network parameters to optimize the computational speed.
Laplace approximation: The idea behind the Laplace approximation is to obtain an approximate posterior around the ’maximum a posterior’ (MAP) estimator of neural network weights with a Gaussian distribution, based on the second derivative of the neural network likelihood functions (MacKay, 1992). The method can be applied post-hoc to a pre-trained neural network model.
@@
Where $\boldsymbol{\hat{\theta}}$ is the MAP estimate and $\mathbf{H}$ is the Hessian matrix. It is the second derivative of the neural network likelihood function regarding the model weights, i.e., $\mathbf{H}_{ij}=-\frac{\partial^{2}}{\partial\boldsymbol{\theta}_{i}\partial\boldsymbol{\theta}_{j}}\log p(\mathbf{Y}|\mathbf{X},\boldsymbol{\theta})$. Fig.  5 (a) illustrates the intuition of Laplace approximation, where the blue density function is the true posterior distribution and the orange Gaussian distribution is the Laplace approximation. This method approximates the posterior locally can simplify the computation of posterior, but the downside is that for the it cannot capture the multi-modal distribution with more than one mode since it is a local estimation around the MAP mode.
@@
Moreover, for DNN, it is still infeasible to compute the invert the Hessian matrix for all parameters, which is typically in the order of several million. Previous work on using Laplace approximation for neural network uncertainty quantification mainly aims to leverage the techniques of Hessian approximation to simplify the computation. One simple solution is to ignore the covariance between weights and only extract the diagonal of the Hessian matrix. Such methods Inspired by the Kronecker factored approximations of the curvature of a neural network, the approach (Ritter et al., 2018b) factorizes the Hessian matrix into the Kronecker product (matrix operation that results in a block matrix) of two smaller matrices. The method brings down the inversion cost of the Hessian matrix and can be scaled to deep convolutional networks and applied to Bayesian online learning efficiently (Ritter et al., 2018a). However, the method introduces an additional assumption that each layer of the neural network is independent (ignoring covariance between layers), which might lead to an overestimation of the variance (uncertainty) in certain directions.
@@
To sum up, the previous methods employed optimization-based schemes like variational inference and Laplace approximations of the posterior. In doing so, strong assumptions and restrictions on the form of the posterior are enforced. The common practice is to approximate the posterior with a Gaussian distribution. The restrictions placed are often credited with inaccuracies induced in predictions and uncertainty quantification performance. The difference between two approximations is that Laplace approximation is a local approximation around the MAP estimation, and it can be applied to a pre-trained neural network and obtain uncertainty quantification without influencing the performance of the neural network. On the other hand, the variational inference is global optimization used during training and may influence the model prediction performance.
Markov Chain Monto Carlo approximation: Markov Chain Monto Carlo (MCMC) is a general method for sampling from an intractable distribution. MCMC constructs an ergodic Markov chain whose stationary distribution is posterior $p(\boldsymbol{\theta}|\mathbf{X},\mathbf{Y})$. Then we can sample from the stationary distribution. The inference step of BNN (Eq. 5) can be approximated as the following equation:
@@
where $\boldsymbol{\theta}_{i}\sim p(\boldsymbol{\theta}|\mathbf{X},\mathbf{Y})$ is a sampled from MCMC. This process generates samples by a Markov chain over a state space, where each sample only depends on the state of the previous sample. The dependency is described with a proposal distribution $T(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta})$ that specifies the probability density of transitioning to a new sample $\boldsymbol{\theta}^{\prime}$ from a given sample $\boldsymbol{\theta}$. Some acceptance criteria are based on the relative density (energy) of two successive samples evaluated at the posterior to determine whether accept the new sample or the previous sample. The vanilla implementation is through the Metropolis-Hastings algorithm (Murphy, 2012) with Gaussian proposal distribution $T(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta})\sim\mathcal{N}(\boldsymbol{\theta},\Sigma)$. Specifically, in each iteration, the algorithm constructs a Markov chain over the state space of $\boldsymbol{\theta}$ with the proposal density distribution. The proposal samples is stochastically accepted with the acceptance probability $\alpha(\boldsymbol{\theta}^{\prime},\boldsymbol{\theta})=\frac{T(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta})p(\boldsymbol{\theta}^{\prime}|x)}{T(\boldsymbol{\theta}|\boldsymbol{\theta}^{\prime})p(\boldsymbol{\theta}|\boldsymbol{x})}$ via a random variable $\mu$ drawn uniformly from the interval $[0,1]$ ($\mu\sim\text{Unif}(0,1)$). If we reject the proposal sample, we retain the previous sample $\boldsymbol{\theta}$. This strategy ensures the stationary distribution of the samples converges to the true posterior after a sufficient number of iterations. However, the isotropic Gaussian proposal distribution shows random walk behavior and can cause slow exploration of the sample space and a high rejection rate. Thus it takes a longer time to converge to the stationary distribution. The problem is more severe because of the high dimensional parameter space of modern DNN and hinders the application of MCMC on neural network parameter sampling (Neal, 2012).
@@
Monto-Carlo (MC) dropout: MC dropout approach (Gal and Ghahramani, 2016) is currently the most popular method for DNN uncertainty quantification in many domains due to its simplicity and ease of implementation. This approach demonstrates that the optimization of a neural network with a dropout layer can be equivalent to approximating a BNN with variational inference on the parametric Bernoulli distribution (Gal and Ghahramani, 2016). Uncertainty estimation can be obtained by computing the variance on multiple stochastic forward predictions with different dropout masks (switching-off some neurons’ activation) as Fig. 5 (b) shows. The average predictions with various weights dropout can be interpreted as doing the approximate integration over the model’s weights (as Eq.5) whose variational distribution follows the Bernoulli distribution. The advantages of MC dropout are: First, it requires little modification to existing DNN architecture design, which allows for straightforward implementation in practice. Second, it mitigates the problem of representing uncertainty by sacrificing prediction accuracy since the methods only influence inference step. However, though there is theoretical justification for the probabilistic interpretation of MC dropout from variational approximation perspective, in many uncertainty benchmark dataset, MC dropout tends to be less calibrated than other baseline UQ methods (Guo et al., 2017).

### v5

@@ 4. Application domains
Biochemistry engineering: Traditional biochemistry discoveries are primarily based on experiments, which are expensive and time-consuming (Bower and Bolouri, 2001). With the rapid advancement of computational science, researchers have turned to simulation and data-driven approaches to aid scientific discovery (Brunton and Kutz, 2022). Simulation methods create mathematical models based on fundamental principles, such as kinetics of microbial growth and product synthesis, fluid dynamics in bioreactors, and design of new proteins and enzymes (Mowbray et al., 2021). However, these models can be computationally intensive and require an in-depth understanding of the biochemical system. In contrast, deep learning-based models use neural networks and a vast amount of biochemistry data to directly extract knowledge and learn complex relationships. However, the common data-driven DNN models are unaware of domain knowledge and face challenges in representing the complexities of biochemical systems, such as the heterogeneity of microbial populations and genetic-environmental variations. This often results in unreliable outputs that require human experts to verify, adding additional labor to the process (Hie et al., 2020). In the realm of drug discovery, although AlphaFold has achieved remarkable success in predicting protein structures—an essential factor in determining protein properties—there remains considerable uncertainty in these predictions. This is because deep learning models lack first-principle biochemical knowledge that would govern such predictions. Thus, Uncertainty-aware deep learning models are crucial as calibrated uncertainty provides information on what they do not know and can potentially help identify novel chemical structures and chemical species whose data may be associated with significant noise (Yang and Li, 2023).
Several factors contribute to the uncertainty of data-driven biochemistry engineering. For example, researchers often have only a partial understanding of the complex biological mechanism and the macro-molecules released from the cells and it is difficult to quantify the heterogeneity in the microbial population, due to genetic and environmental variations. For different tasks, these factors may play a different role in the UQ for the neural network model. In protein engineering, the protein function prediction (based on protein sequences, protein structures, and protein-protein interactions) is challenging because of the large amount of diversity of protein folds and the lack of a complete understanding of the protein structure (Bradford et al., 2018). For bioreactor engineering, the prediction of bioreactor system performance is affected by the intrinsic interactions between bioreactor operating factors (e.g., pH, temperature, and substrate concentration). An iterative method has been used to enable the propagation of model uncertainty for multi-step ahead predictions (Mowbray et al., 2021).
Natural Language Processing: The advancement of pre-trained language models (PLMs) has revolutionized language processing by addressing various tasks in a unified manner (Li et al., 2021). Although natural language processing (NLP) has made remarkable strides with the emergence of large language models (LLMs), these models are prone to hallucination, often generating misleading or fabricated content (Manakul et al., 2023a; Xu et al., 2024). Thus, uncertainty quantification plays a critical role in improving the trustworthiness of LLMs in several applications, such as assessing the reliability of machine translations, gauging sentiment polarity in sentiment analysis, especially in ambiguous cases, and improving the accuracy and reliability of speech-to-text models in speech recognition. However, quantifying uncertainty in LLMs presents significant challenges. First, uncertainty can arise from various sources related to both data and model perspectives. Data uncertainty stems from ambiguities, noise (such as out-of-vocabulary words or distractions), and semantic complexities in the input language. Model uncertainty occurs when the model lacks the specific knowledge required for out-of-distribution input queries, leading to arbitrary responses, a phenomenon known as the hallucination problem in LLMs. Second, the large vocabulary space complicates the direct assessment of confidence through probability likelihood, and relying solely on prediction logits can lead to overconfidence (Guo et al., 2017). Moreover, capturing uncertainty in the semantic space is essential due to the semantic similarity among different answers. Third, aligning uncertainty with actual correctness (calibration) is crucial for trustworthy applications of large language models.
Engineering Design: Engineering simulation accelerates the development of reliable and repeatable design processes. However, the computing resource consumption is dramatically raised in the whole development process. Deep learning-based methods have great potential to accelerate modern industrial product design in several aspects. First, the deep learning-based surrogate model for the physics process has the potential to accelerate the traditional simulation-based engineering process, for example, in material modeling (Liu and Wang, 2019), flow sculpting (Lee et al., 2019), distributed energy resources optimization (Odonkor and Lewis, 2019). Second, deep generative models have been successfully utilized for synthesis. For example, generative adversarial networks are leveraged to synthesize hierarchical designs with inter-part dependencies (Chen and Fuge, 2019), intelligent computer-aided design (CAD) (Chen and Fuge, 2019), 2D wheel design (Oh et al., 2019). Third, leveraging deep learning models can extract human preferences and design strategies from a large number of observation data (Suryadi and Kim, 2019). However, in these applications, a major concern for deep learning in engineering design is the lack of trustworthiness in real-world applications. Through explicit uncertainty quantification, the model is aware of the potential mistakes it can make and facilitates broader adoption of deep learning in high-stakes decision-making, including manufacturing, aviation, and automotive engineering. The uncertainty source comes from diverse aspects, including data uncertainty due to measurement noise and randomness in the physics variable. Second is model uncertainty due to limited data availability in the engineering domain.
@@ 5. A Taxonomy of DNN UQ Methodology
In this section, we provide a new taxonomy of existing DNN UQ methods based on the type of uncertainty sources each method can capture. As mentioned before, two main sources of uncertainty exist in the DNN predictions, and for different domain applications, different sources may play a major role in affecting the model prediction. Thus it is vital to understand which type of uncertainty source each method can address. In general, we categorize existing works into three branches based on the uncertainty sources as Fig. 5 shows: model uncertainty, data uncertainty, and the combination of the two. We briefly introduce each category as follows:
  • Model uncertainty: In this category, the approaches consider the model uncertainty resulting from the parameters (BNN), architectures (ensemble methods), or sample distribution (deep Gaussian process and distance-aware model).
  • Data uncertainty: The approaches in this category aim to account for the uncertainty from the inherent randomness or noise in the data. The general idea is to construct a distribution over the prediction. These approaches are further split into deep discriminative models and deep generative models.
  • The combination of model and data uncertainty: Many approaches aim to capture the total uncertainty in one framework, which is the combination of model and data uncertainty. The straightforward approach combines the approaches in data and model uncertainty, forming a coherent framework but introducing more computation and storage demand. Another framework called evidential deep learning overcomes the computational challenge with a single neural network model.
@@ 5.1. Model Uncertainty
This subsection reviews the existing framework for DNN model uncertainty. We categorize the approaches into three subcategories: Bayesian neural network, ensemble models, and sample distribution-related models. We will introduce each subcategory in detail.
@@ 5.1.1. Bayesian Neural Networks
From a frequentist point of view, there exists a single set of parameters $\boldsymbol{\boldsymbol{\theta}}*$ that fit the DNN model best, where $\boldsymbol{\theta}*=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathcal{L}(\mathbf{Y},f(\mathbf{X},\boldsymbol{\theta}))$ and $\mathcal{L}$ is the loss function. However, the point estimation of DNN parameters tends to be over-fitting and overconfident (Thulasidasan et al., 2019). In contrast, the Bayesian neural network (BNN) imposes a prior on the neural network parameters $p(\boldsymbol{\theta})$ and learns the posterior distribution of these parameters $p(\boldsymbol{\theta}|\mathbf{X},\mathbf{Y})$.
@@
For inference on the new samples $\boldsymbol{x}*$, we can marginalize out the model parameters by:
@@
However, the approach is analytically intractable and does not have a closed-form solution and an approximation must be made for prediction in BNN. To estimate the posterior of the neural network parameters, various approaches have been proposed to approximate the parameter posterior in a simpler form to solve it in a tractable way. Some approximation methods define a parameterized class of distributions, $\mathcal{Q}$, from which they select an approximation $q_{\phi}(\boldsymbol{\theta})$for the posterior. For example, $\mathcal{Q}$ can be the set of all factorized Gaussian distributions, and $\phi$ is the parameters of mean and diagonal variance. The distribution $q_{\phi}(\boldsymbol{\theta})\in\mathcal{Q}$ is selected according to some optimization criteria to approximate the posterior. Two popular approaches for optimization are variational inference (Blei et al., 2017) and Laplace approximation (Friston et al., 2007). Instead of approximating the posterior analytically, another approach aims to solve this problem by Monte Carlo sampling called Markov Chain Monte Carlo sampling. In the following, we will review existing methodologies in the three subcategories in detail.
@@
One challenge of variational inference lies in the choice of the parameterized class of distribution $q_{\boldsymbol{\phi}}(\boldsymbol{\theta})$. The original method relies on the use of Gaussian approximating distribution with a diagonal covariance matrix (mean-field variational inference) (Hinton and Van Camp, 1993; Posch et al., 2019). This method leads to a straightforward lower bound for optimization, but the approximation capability is limiting. To capture the posterior correlations between parameters, the second category of approaches extends the diagonal covariance to general covariance matrix while still leading to a tractable algorithm by maximizing the above ELBO (Posch and Pilz, 2020). However, the full covariance matrix not only increases the number of trainable parameters but also introduces substantial memory and computational cost for DNN. To reduce the computation, the third category of approaches aims to simplify the covariance matrix structure with certain assumptions. Some approaches assume independence among layers, resulting in block-diagonal structure covariance matrix (Sun et al., 2017; Zhang et al., 2018). Others find that for a variety of deep BNN trained using Gaussian variational inference, the posterior consistently exhibits strong low-rank structure after convergence (Swiatkowski et al., 2020), they propose to decompose the dense covariance matrix into a low-rank factorization to simplify the computation. Additionally, (Mishkin et al., 2018) assumes the covariance matrix taking the form of “diagonal plus low-rank" structure for a more flexible and faster approximation. Another line of work introduces sparse uncertainty structure via a hierarchical posterior or employing normalizing flow with low-dimensional auxiliary variables (Ritter et al., 2021; Louizos and Welling, 2017) to reduce the computation.
The disadvantage of variational Gaussian approximation for DNN parameters is that to capture the full correlations among model latent weights, it requires a large number of variational parameters to be optimized, which scales quadratically with the number of latent weights in the model. Existing works aim to simplify the covariance structure with certain assumptions while still capturing the correlation between neural network parameters to optimize the computational speed.
Laplace approximation: The idea behind the Laplace approximation is to obtain an approximate posterior around the ’maximum a posterior’ (MAP) estimator of neural network weights with a Gaussian distribution, based on the second derivative of the neural network likelihood functions (MacKay, 1992). The method can be applied post hoc to a pre-trained neural network model.
@@
Where $\boldsymbol{\hat{\theta}}$ is the MAP estimate and $\mathbf{H}$ is the Hessian matrix. It is the second derivative of the neural network likelihood function regarding the model weights, i.e., $\mathbf{H}_{ij}=-\frac{\partial^{2}}{\partial\boldsymbol{\theta}_{i}\partial\boldsymbol{\theta}_{j}}\log p(\mathbf{Y}|\mathbf{X},\boldsymbol{\theta})$. Fig.  6 (a) illustrates the intuition of Laplace approximation, where the blue density function is the true posterior distribution and the orange Gaussian distribution is the Laplace approximation. This method approximates the posterior locally and can simplify the computation of the posterior, but the downside is that it cannot capture the multi-modal distribution with more than one mode since it is a local estimation around the MAP mode.
@@
Moreover, for DNN, it is still infeasible to compute the inverted Hessian matrix for all parameters, which is typically in the order of several million. Previous work on using Laplace approximation for neural network uncertainty quantification mainly aims to leverage the techniques of Hessian approximation to simplify the computation. One simple solution is to ignore the covariance between weights and only extract the diagonal of the Hessian matrix. Such methods Inspired by the Kronecker factored approximations of the curvature of a neural network, the approach (Ritter et al., 2018b) factorizes the Hessian matrix into the Kronecker product (matrix operation that results in a block matrix) of two smaller matrices. The method brings down the inversion cost of the Hessian matrix and can be scaled to deep convolutional networks and applied to Bayesian online learning efficiently (Ritter et al., 2018a). However, the method introduces an additional assumption that each layer of the neural network is independent (ignoring covariance between layers), which might lead to an overestimation of the variance (uncertainty) in certain directions.
@@
To sum up, the previous methods employed optimization-based schemes like variational inference and Laplace approximations of the posterior. In doing so, strong assumptions and restrictions on the form of the posterior are enforced. The common practice is to approximate the posterior with a Gaussian distribution. The restrictions placed are often credited with inaccuracies induced in predictions and uncertainty quantification performance. The difference between the two approximations is that the Laplace approximation is a local approximation around the MAP estimation, and it can be applied to a pre-trained neural network and obtain uncertainty quantification without influencing the performance of the neural network. On the other hand, the variational inference is a global optimization used during training and may influence the model prediction performance.
Markov Chain Monte Carlo approximation: Markov Chain Monte Carlo (MCMC) is a general method for sampling from an intractable distribution. MCMC constructs an ergodic Markov chain whose stationary distribution is posterior $p(\boldsymbol{\theta}|\mathbf{X},\mathbf{Y})$. Then we can sample from the stationary distribution. The inference step of BNN (Eq. 4) can be approximated as the following equation:
@@
where $\boldsymbol{\theta}_{i}\sim p(\boldsymbol{\theta}|\mathbf{X},\mathbf{Y})$ is a sampled from MCMC. This process generates samples by a Markov chain over a state space, where each sample only depends on the state of the previous sample. The dependency is described with a proposal distribution $T(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta})$ that specifies the probability density of transitioning to a new sample $\boldsymbol{\theta}^{\prime}$ from a given sample $\boldsymbol{\theta}$. Some acceptance criteria are based on the relative density (energy) of two successive samples evaluated at the posterior to determine whether to accept the new sample or the previous sample. The vanilla implementation is through the Metropolis-Hastings algorithm (Murphy, 2012) with Gaussian proposal distribution $T(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta})\sim\mathcal{N}(\boldsymbol{\theta},\Sigma)$. Specifically, in each iteration, the algorithm constructs a Markov chain over the state space of $\boldsymbol{\theta}$ with the proposal density distribution. The proposal samples is stochastically accepted with the acceptance probability $\alpha(\boldsymbol{\theta}^{\prime},\boldsymbol{\theta})=\frac{T(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta})p(\boldsymbol{\theta}^{\prime}|x)}{T(\boldsymbol{\theta}|\boldsymbol{\theta}^{\prime})p(\boldsymbol{\theta}|\boldsymbol{x})}$ via a random variable $\mu$ drawn uniformly from the interval $[0,1]$ ($\mu\sim\text{Unif}(0,1)$). If we reject the proposal sample, we retain the previous sample $\boldsymbol{\theta}$. This strategy ensures the stationary distribution of the samples converges to the true posterior after a sufficient number of iterations. However, the isotropic Gaussian proposal distribution shows random walk behavior and can cause slow exploration of the sample space and a high rejection rate. Thus it takes a longer time to converge to the stationary distribution. The problem is more severe because of the high dimensional parameter space of modern DNN which hinders the application of MCMC on neural network parameter sampling (Neal, 2012).
@@
Monte-Carlo (MC) dropout: MC dropout approach (Gal and Ghahramani, 2016) is currently the most popular method for DNN uncertainty quantification in many domains due to its simplicity and ease of implementation. This approach demonstrates that the optimization of a neural network with a dropout layer can be equivalent to approximating a BNN with variational inference on the parametric Bernoulli distribution (Gal and Ghahramani, 2016). Uncertainty estimation can be obtained by computing the variance on multiple stochastic forward predictions with different dropout masks (switching off some neurons’ activation) as Fig. 6 (b) shows. The average predictions with various weights dropout can be interpreted as doing the approximate integration over the model’s weights (as Eq.4) whose variational distribution follows the Bernoulli distribution. The advantages of MC dropout are: First, it requires little modification to existing DNN architecture design, which allows for straightforward implementation in practice. Second, it mitigates the problem of representing uncertainty by sacrificing prediction accuracy since the methods only influence the inference step. However, though there is theoretical justification for the probabilistic interpretation of MC dropout from a variational approximation perspective, in many uncertainty benchmark datasets, MC dropout tends to be less calibrated than other baseline UQ methods (Guo et al., 2017).